Freddie gets a car ride to the top of a hill, then rides his skateboard back down. The downhill ride takes 15 minutes longer than the uphill. If the car averages 32 mph, and Freddie rides his skateboard at 8 mph, find out how much time he spends skateboarding.
Let t=time spent driving to the top of the hill then, t+15=time going down the hill.
distance up=distance down
from the formula speed=distance/time(s=d/t), then d=st,
32t=8(t+15)
32t=8t+120
I can't get to the answer of 1/3 hour. I keep coming up with 1/4 hour.

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

15 minutes is 1/4 hr so t+1/4 not t+15
32t=8(t+1/4)
32t=8t+2
24t=2
t=1/12 hour.
t=5 minutes
5+15=20 minutes
20/60=1/3 hour skateboarding
If you use t+15 then you get t=5 but 5 means 5 hours not 5 minutes!!!

Answer by stanbon(75887)

Freddie gets a car ride to the top of a hill, then rides his skateboard back down.
The downhill ride takes 15 minutes longer than the uphill.
If the car averages 32 mph, and Freddie rides his skateboard at 8 mph, find out how much time he spends skateboarding.
---
Drive up DATA:
rate = 32 mph ; time = t hrs ; distance = 32t miles
-----
Skate down DATA:
rate = 8 mph ; time = t+(1/4) hrs ; distance = 8(t+(1/4)) miles
-----
Equation:
distance up = distance down:
32t = 8t+2
24t = 2
time = 1/12 hr = 5 minutes (time up)
time down = 5 min + 15 min = 20 min (time down)
---
Total skate time = 20 min = 1/3 hr.
====================================
Cheers,
Stan H.
===================